OpenStudy (anonymous):

Graph the equation by plotting points. Give the coordinates of the vertex. y=-2x^2+4

OpenStudy (anonymous):

Find the two zeros of the function. They are +/- square root of 2 (you can get this by doing: 0=-2x^2+4=-2(x^2-2)=-2(x-sqrt2)*(x+sqrt2). As this is a simple quadratic function, the x coordinate of the vertex will be the in the middle of the two zeros, which happens to be 0. You insert 0 into the equation and get 4, the vertex therefore has coordinates (0, 4). Plot the graph through these 3 points (the two zeros and the vertex).